OpenStudy (anonymous):

what is the solution of this equation? 1/2x=2

OpenStudy (anonymous):

x=4 (I think you mean this (1/2)x=2

OpenStudy (jamesj):

Or if 1/(2x) = 2, then 4x = 1 => x = 1/4